# Topology of ultra-localized insulators and superconductors

### v1

@@ Abstract
The topology of an insulator can be defined even when all eigenstates of the system are localized — an extreme case of Anderson insulators that we call ultra-localized. We derive the classification of such ultra-localized insulators in all symmetry classes and dimensions. We clarify their bulk-boundary correspondence and show that ultra-localized systems are in many instances phases of matter not described by the known classification of topological insulators and superconductors.
@@
Band insulators are thus only insulating at very specific electron filling, while Anderson insulators can be insulating over entire ranges of fillings – depending on the disorder strength, symmetry and dimension of the system Evers and Mirlin (2008). At very strong disorder, Anderson insulators are insulating over their whole spectrum. We will refer to such fully insulating systems as “ultra-localized insulators”, or “ultra-insulators” (UI) for short. While in one-dimension all Anderson insulators are ultra-insulators, in higher dimensions this requires a large but finite disorder. Ultra-insulators are non-conducting even at arbitrary temperature or in non-equilibrium settings such as under strong applied voltages. Thus, the notion of ultra-insulator is naturally relevant to periodically driven systems Rudner et al. (2013); Titum et al. (2016); Nathan et al. (2017), where stabilizing insulating phases is often a challenging task.
@@
Figure 1: Localization properties of eigenstates of generic disordered systems across topological phase transitions (a) Phase transition between a topological (ultra localized) insulator and a trivial insulator in $d=1$, driven by some parameter $\lambda$. In both phases all states are localized (grey), and critical states (blue) appear at zero energy Essin and Gurarie (2015); Bagrets et al. (2016) at the transition point. (b) Phase transition out of a disordered TI (hatched region) in $d=2$, not deformable into an ultra-insulator. The states in red are metallic. A trivial ultra-insulator emerges at strong disorder as the delocalized states pair-annihilate, while TIs appear only at specific fillings and require bulk delocalized states. Here, $W$ is the disorder strength. (c) Transition between a topological and a trivial ultra-insulator in $d=3$. There is an extended critical region characterized by delocalized states appearing somewhere in the bulk spectrum.
@@
Classification — We consider a $d$-dimensional TUI on a lattice $\mathcal{L}$ with sites $\mathbf{r}\in\mathcal{L}$ in $d$-dimensional Euclidean space, assuming one orbital per site for simplicity. By definition, the system is characterized by a fully-localized bulk spectrum. It follows that its eigenstates $|\psi_{\mathbf{n}}\rangle$ can be indexed by $d$-dimensional vectors $\mathbf{n}\in\mathcal{L}$, that can always be chosen such that the eigenstate $|\psi_{\mathbf{n}}\rangle$ (at energy $E_{\mathbf{n}}$) is localized near site $\mathbf{r}=\mathbf{n}$, although such an assignment is not unique. The eigenstates $|\psi_{\mathbf{n}}\rangle$ can be seen as the columns of a unitary matrix $U_{\mathbf{r},\mathbf{n}}$,
@@
With the help of the unitary matrix $U$, one may construct the auxiliary chiral Hamiltonian
@@
We use the mapping to the auxiliary Hamiltonian (3) to classify TUIs in all dimensions and classes. The complete classification is summarized in Table 1, and details on its derivation for each class are discussed in SM . Crucially, while the topology of TUIs is invariant under a local relabelling of the eigenstates (i.e., a relabelling that leaves the unitary $U$ local), this may change the topology of the auxiliary Hamiltonian (3) in $d=1$, but this relabelling freedom is inconsequential in higher dimensions SM . As a result, the period-eight Bott periodicity of the TUI classification sets in for $d\geq 2$ only, whereas $d=1$ is an exceptional case that never repeats in higher dimensions. The Bott clock, which relates topological phases of different classes and different dimensions, is absent in our classification. Instead, Table 1 has a threefold structure, where the resulting classification is associated to one of the three chiral classes AIII, BDI, and CII.
class $\,d=1$ $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
AIII $\mathbb{Z}$ 00 $\mathbb{Z}\times\mathbb{Z}$ 00 $\mathbb{Z}\times\mathbb{Z}$ 00 $\mathbb{Z}\times\mathbb{Z}$ 00 $\mathbb{Z}\times\mathbb{Z}$
A 00 00 $\mathbb{Z}$ 00 $\mathbb{Z}$ 00 $\mathbb{Z}$ 00 $\mathbb{Z}$
DIII $\mathbb{Z}_{2}$ 00 $\mathbb{Z}$ 00 $\mathbb{Z}$ 00 $\mathbb{Z}$ 00 $\mathbb{Z}$
CI 00 00 $\mathbb{Z}$ 00 $\mathbb{Z}$ 00 $\mathbb{Z}$ 00 $\mathbb{Z}$
BDI $\mathbb{Z}$ 00 00 00 $2\mathbb{Z}\times 2\mathbb{Z}$ 00 $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$$\mathbb{Z}\times\mathbb{Z}$
[…]
CII $2\mathbb{Z}$ 00 $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$$\mathbb{Z}\times\mathbb{Z}$ 00 00 00 $2\mathbb{Z}\times 2\mathbb{Z}$
[…]
Table 1: Classification of topological ultra-insulators and superconductors. We highlight a threefold structure of the classification (with an exception in $d=1$), with three groups associated to the three chiral classes. The bulk topological invariants associated to each of these systems are summarized in SM . “$2\mathbb{Z}$” indicates that the invariant takes even values only.
@@
Wannier localizability of TIs — Certain topological phases, such as quantum (spin) Hall insulators and three-dimensional TIs, have an obstruction to Wannier localization of their bulk spectrum Chalker and Coddington (1988); Onoda et al. (2007a); Morimoto et al. (2015), and thus cannot be deformed to TUIs without closing the mobility gap. Others, such as the one-dimensional topological phases, can be adiabatically connected to an ultra-localized limit Kohn (1959). It is then natural to ask which topological phases of the conventional TI classification can exist as ultra-localized insulators and superconductors Monaco et al. (2016); Panati (2018); Marcelli et al. (2023). Since for $d>0$ topological equivalence as TUI phases implies topological equivalence according to the TI classification, it is possible to assign a TI invariant to each TUI. The result of such an assignment is shown in Table 2: A $\checkmark$” in Tab. 2 indicates that all topological phases in the corresponding symmetry class exist as TUI (i.e., that the TI invariants calculated for TUIs span the full set of allowed TI invariants for that symmetry class), whereas a “00” in Tab. 2 indicates that TUIs have trivial TI invariant.
For the Wigner-Dyson classes A, AI, AII, which do not have a “special” energy $E=0$, the energy eigenvalues of a TUI may be chosen such that all eigenstates have negative energy. The resulting conventional invariants are always trivial, as indicated in Tab. 2. The opposite is true for chiral classes AIII, BDI and CII, where all conventional topological phases can be localized. As a consequence, the topology of, e.g., chiral TIs in $d=3$ does not imply the existence of delocalized bulk states Altland et al. (2024). Whether the superconducting classes C, CI, D, and DIII are Wannier localizable or not depends on the dimension and the symmetry class, as summarized in Table 2 (see SM for details).
While topological equivalence as TUIs implies topological equivalence as conventional TI if $d>0$, the opposite is not true. In this sense, the TUI classification is more “refined” than the TI classification. An nontrivial example is the $\mathbb{Z}\times\mathbb{Z}$ classification of chiral TUIs with broken time-reversal symmetry for $d=3$ (see Table 1), which has to be contrasted with the $\mathbb{Z}$ classification of TIs: For each (integer-valued) conventional topological invariant, an additional integer-valued topological invariant is required to characterize the TUI. This additional invariant counts the number of finite-energy critical states appearing on the boundary SM ; Altland et al. (2024). Table 4 of SM gives an overview of this “beyond-tenfold-way” aspect of the TUI classification for all symmetry classes and all dimensions.
class $\,d=1$ $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
A 00 00 00 00 00 00 00 00 00
AIII $\checkmark$ 00 $\checkmark$ 00 $\checkmark$ 00 $\checkmark$ 00 $\checkmark$
AI 00 00 00 00 00 00 00 00 00
BDI $\checkmark$ 00 00 00 $\checkmark$ 00 $\checkmark$ $\checkmark$ $\checkmark$
D $\checkmark$ 00 00 00 00 00 00 $\checkmark$ $\checkmark$
DIII $\checkmark$ 00 $0/\checkmark$ 00 00 00 $\checkmark$ 00 $\checkmark$
AII 00 00 00 00 00 00 00 00 00
CII $\checkmark$ 00 $\checkmark$ $\checkmark$ $\checkmark$ 00 00 00 $\checkmark$
C 00 00 00 $\checkmark$ $\checkmark$ 00 00 00 00
CI 00 00 $\checkmark$ 00 $\checkmark$ 00 $0/\checkmark$ 00 00
@@ D $\checkmark$ 00 00 00 00 00 00 $\checkmark$ $\checkmark$
Conclusions — Ultra-localized insulators, insulators that have exponentially localized states at all energies, can nevertheless be topological. The resulting topology is a global feature of the fully localized spectrum, and only transitions between topologically inequivalent ultra-insulators lead to metallic states somewhere in the spectrum. While these systems coincide with conventional topological matter in one dimension, and cannot exist in two dimensions, we uncovered many new instances of TUIs in three dimensions in different symmetry classes. As a hallmark their topology, TUIs host anomalous delocalized surface states that resist arbitrary surface disorder.
@@
The fact that TUI and TI classifications differ, is a consequence of the different objectives – studying equivalence classes of fully localized Hamiltonians versus those of ground states of gapped Hamiltonians. This change of viewpoint may be fruitful when applied to time-dependent quantum systems, such as Floquet and random unitary circuits, where local unitary operators (instead of Hamiltonians) play a central role Titum et al. (2016); Roy and Harper (2017); Long et al. (2021); Liu et al. (2023b); Fidkowski and Hastings (2023). It would then be desirable to extend our classification scheme of local unitaries Zhang and Levin (2021) beyond the single particle description.
@@
The construction of TUIs presented in this work corresponds to an infinite-disorder limit of a particular correlated disorder, instead of commonly considered independently distributed disorder. It is an open question whether some TUIs can be reached by adding generic (uncorrelated) disorder in a clean topological insulator. Since the limit of infinite uncorrelated disorder gives rise to a trivial ultra-insulator, a competition between different disorders should result in a transition between TUIs, as sketched in Fig. 1(c). More extreme types of structural disorders, such as amorphous crystals, have recently been shown to realize conventional topological insulators Agarwala and Shenoy (2017); Mitchell et al. (2018); Marsal et al. (2020), but the full topological classification of disordered electronic matter is unknown, in particular extensions to disordered hyperuniform systems Torquato and Stillinger (2003); Zheng et al. (2020). Our work thus opens a pathway to generalize notions of topology to wider range of systems.
@@ Acknowledgements.
Acknowledgments — B.L. acknowledges financial support from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 214461). L.T. acknowledges support by the FNS/SNF Ambizione Grant No. PZ00P2_179962. T.N. acknowledges support from the Swiss National Science Foundation through a Consolidator Grant (iTQC, TMCG-$2\_213805$) P.B. is supported by the Deutsche Forschungsgemeinschaft (DFG) project Grant No. 277101999 within the CRC network TR 183 (subproject No. A03).
@@
To find $K_{\rm aux}^{\prime}(d)$, we note that the Hamiltonian $H$ of the UI determines the eigenstates $|\psi_{\mathbf{n}}\rangle$ up to a phase factor $e^{i\phi_{\mathbf{n}}}$ and an arbitrary relabeling of the indices $\mathbf{n}$, as long as such a relabelling preserves the locality of $U$. The uncertainty of the phase factors means that the matrix $U$ may be right-multiplied by a diagonal matrix $D$ containing phase factors $e^{i\phi_{\mathbf{n}}}$. The relabeling of the indices $\mathbf{n}$ amounts to the degree of freedom to right-multiply $U$ by a “relabeling matrix” $P_{\mathbf{n}^{\prime},\mathbf{n}}$. The relabeling matrix is a permutation matrix, i.e., its entries are 00 or $1$. Because of the condition that the site indices $\mathbf{n}$ are chosen such that the eigenstate $|\psi_{\mathbf{n}}\rangle$ is localized near lattice position $\mathbf{n}$, the relabeling matrix is local in the site indices $\mathbf{n}$ and $\mathbf{n}^{\prime}$, i.e., $P_{\mathbf{n}^{\prime},\mathbf{n}}=0$ if $|\mathbf{n}-\mathbf{n}^{\prime}|$ exceeds the localization length. Auxiliary Hamiltonians that are obtained from the trivial reference Hamiltonian $H_{\mathrm{aux}}^{0}$ — for which we take the eigenvector matrix $𝟙U^{0}=\openone$ — by permutation of the eigenstate labels or by changing the phases of the eigenstates ${|\psi_{\mathbf{n}}\rangle}$ are of the form
@@
Correspondingly, the matrix $U$ is an orthogonal matrix and the auxiliary Hamiltonian $H_{\mathrm{aux}}$ is in tenfold-way class BDI, so that $K_{\rm aux}(d)={\cal K}_{\rm BDI}(d)$ for class AI. In this case, the “phase matrix” $D$ consists of signs ±$\pm 1$. Since relabeling matrices $P$ may have arbitrary integer winding number in $d=1$, but possess no nontrivial topology if $d>1$, it also follows that $K_{\rm aux}^{\prime}(d)={\cal K}_{\rm BDI}(1)=\mathbb{Z}$ if $d=1$ and $K_{\rm aux}^{\prime}(d)=0$ otherwise. We conclude that
@@
Class AII.— For class AII, $H$ must satisfy Eq. (4). Eigenvalues of $H$ come in twofold (Kramers) degenerate pairs. The same symmetry condition applies to the eigenvector matrix $U$, which makes it a symplectic matrix, and to the auxiliary Hamiltonian $H_{\mathrm{aux}}$, which places it in tenfold-way class CII. Since the relabeling matrix $P$ is proportional to the $2\times 2$ unit matrix $\sigma_{0}$, the corresponding winding numbers for $d=1$ are always even, so that $K_{\rm aux}^{\prime}(d)={\cal K}_{\rm CII}(1)=2\mathbb{Z}$ if $d=1$ and 00 otherwise. We then find
@@
where $\tau_{3}$ is the third Pauli matrix in a suitably chosen local basis. Further, in classes BDI and CII, $H$ is subject to an additional symmetry condition of the form $H=H^{*}$ or $H=\sigma_{2}H^{*}\sigma_{2}$. This additional symmetry condition commutes with $\tau_{3}$ and imposes that elements of $H$ are real or quaternion numbers. The antisymmetry condition (S8) also imposes that the eigenvalues come in pairs ±$\pm E_{\mathbf{n}}$ and that the corresponding eigenvectors are related to each other by multiplication by $\tau_{3}$. In class CII, all eigenvalues are twofold (Kramers) degenerate.
@@
[…]
@@
Its eigenvalues come in pairs ±$\pm E_{\mathbf{n}}$ and the eigenvectors are related to each other via particle-hole conjugation (i.e., multiplication with $\tau_{1}$ and complex conjugation). For an ultra-localized Hamiltonian we impose the additional condition that this also applies to the eigenvalue $E=0$, i.e., we rule out the possibility of a “topological zero-eigenvalue” for which the eigenstates are localized, but not related by particle-hole conjugation.
@@
where $\Lambda$ is a diagonal matrix containing blocks $E_{\mathbf{n}}\tau_{3}$, with each pair ±$\pm E_{\mathbf{n}}$ of eigenvalues of $H$ appearing once. The eigenvector matrix $U$ satisfies the antiunitary symmetry

### v2

@@ Abstract
The topology of an insulator can be defined even when all eigenstates of the system are localized — an extreme case of Anderson insulators that we call ultra-localized. We derive the classification of such ultra-localized insulators in all symmetry classes and dimensions. We clarify their bulk-boundary correspondence and show that ultra-localized systems are in many instances phases of matter not described by the known classification of topological insulators and superconductors. As a consequence, we clarify which conventional topological phases are Wannierizable, and which topological phases cannot exist without delocalized states.
@@
Band insulators are thus only insulating at very specific electron filling, while Anderson insulators can be insulating over entire ranges of fillings – depending on the disorder strength, symmetry and dimension of the system Evers and Mirlin (2008). At very strong disorder, Anderson insulators are insulating over their whole spectrum. We will refer to such fully insulating systems as “ultra-localized insulators”, or “ultra-insulators” (UI) for short. While in one-dimension all Anderson insulators are ultra-insulators, in higher dimensions this requires a large but finite disorder strength. Ultra-insulators are non-conducting even at arbitrary temperature or in non-equilibrium settings such as under strong applied voltages. Thus, the notion of ultra-insulator is naturally relevant to periodically driven systems Rudner et al. (2013); Titum et al. (2016); Nathan et al. (2017), where stabilizing insulating phases is often a challenging task.
@@
Figure 1: Localization properties of eigenstates of generic disordered systems across topological phase transitions. (a) Phase transition between a topological (ultra localized) insulator and a trivial insulator in $d=1$, driven by a parameter $\lambda$. In both phases all states are localized (grey), and critical states (blue) appear at zero energy Essin and Gurarie (2015); Bagrets et al. (2016) at the transition point. (b) Phase transition out of a disordered TI (hatched region) in $d=2$, not deformable into an ultra-insulator. The states in red are metallic. A trivial ultra-insulator emerges at strong disorder as the delocalized states pair-annihilate, while TIs appear only at specific fillings and require bulk delocalized states. Here, $W$ is the disorder strength. (c) Transition between a topological and a trivial ultra-insulator in $d=3$. There is an extended critical region characterized by delocalized states appearing somewhere in the bulk spectrum.
@@
Classification — We consider a $d$-dimensional TUI on a lattice $\mathcal{L}$ with sites $\mathbf{r}\in\mathcal{L}$ in $d$-dimensional Euclidean space, assuming one orbital per site for simplicity. By definition, the system is characterized by a localized bulk spectrum. It follows that its eigenstates $|\psi_{\mathbf{n}}\rangle$ can be indexed by $d$-dimensional vectors $\mathbf{n}\in\mathcal{L}$, that can always be chosen such that the eigenstate $|\psi_{\mathbf{n}}\rangle$ (at energy $E_{\mathbf{n}}$) is localized near site $\mathbf{r}=\mathbf{n}$, although such an assignment is not unique. The eigenstates $|\psi_{\mathbf{n}}\rangle$ can be seen as the columns of a unitary matrix $U_{\mathbf{r},\mathbf{n}}$,
@@
With the help of the unitary matrix $U$, one may construct the auxiliary chiral Hamiltonian Roy and Harper (2017)
@@
We use the mapping to the auxiliary Hamiltonian (3) to classify TUIs in all dimensions and classes. The complete classification is summarized in Table 1, and details of its derivation for each class are discussed in SM . Crucially, while the topology of TUIs is invariant under a local relabelling of the eigenstates (i.e., a relabelling that leaves the unitary $U$ local), this may change the topology of the auxiliary Hamiltonian (3) in $d=1$, but this relabelling freedom is inconsequential in higher dimensions SM . As a result, the period-eight Bott periodicity of the TUI classification sets in for $d\geq 2$ only, whereas $d=1$ is an exceptional case that never repeats in higher dimensions. The Bott clock, which relates topological phases of different classes and different dimensions, is absent in our classification. Instead, Table 1 has a threefold structure, where the resulting classification is associated to one of the three chiral classes AIII, BDI, and CII. We note that for $d>1$ our classification table conicides with the topological classification of a totally different systems, i.e., that of point-gap non-Hermitian systems Gong et al. (2018).
class d=1\,d=1 $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
AIII $\mathbb{Z}$ 0 $\mathbb{Z}\times\mathbb{Z}$ 0 $\mathbb{Z}\times\mathbb{Z}$ 0 $\mathbb{Z}\times\mathbb{Z}$ 0 $\mathbb{Z}\times\mathbb{Z}$
A 0 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$
DIII $\mathbb{Z}_{2}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$
CI 0 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$
BDI $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}\times 2\mathbb{Z}$ 0 $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $\mathbb{Z}\times\mathbb{Z}$
AI 0 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$
D $\mathbb{Z}_{2}$ 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$
CII $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $\mathbb{Z}\times\mathbb{Z}$ 0 0 0 $2\mathbb{Z}\times 2\mathbb{Z}$
AII 0 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$
C 0 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$
Table 1: Classification of topological ultra-insulators and superconductors. The classification has a threefold structure (with an exception in $d=1$), with three groups associated to the three chiral classes. The bulk topological invariants associated to each of these systems are summarized in SM . $2\mathbb{Z}$” indicates that the invariant takes even values only.
@@ D $\mathbb{Z}_{2}$ 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$
Wannier localizability of TIs — Our classification of TUIs can be used to determine an important property of conventional topological insulator phases. Certain topological phases, such as quantum (spin) Hall insulators and three-dimensional TIs, have an obstruction to Wannier localization of their bulk spectrum Chalker and Coddington (1988); Onoda et al. (2007a); Morimoto et al. (2015), and thus cannot be deformed to TUIs without closing the mobility gap. Others, such as the one-dimensional topological phases, can be adiabatically connected to an ultra-localized limit Kohn (1959). It is then natural to ask which topological phases of the conventional TI classification can exist as ultra-localized insulators and superconductors Monaco et al. (2016); Panati (2018); Marcelli et al. (2023). Since for $d>0$ topological equivalence as TUI phases implies topological equivalence according to the TI classification, it is possible to assign a TI invariant to each TUI. The result of such an assignment is shown in Table 2: A $\checkmark$” in Tab. 2 indicates that all topological phases in the corresponding symmetry class exist as TUI (i.e., that the TI invariants calculated for TUIs span the full set of allowed TI invariants for that symmetry class), whereas a “0” in Tab. 2 indicates that TUIs have trivial TI invariant.
class d=1\,d=1 $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
A 0 0 0 0 0 0 0 0 0
AIII $\checkmark$ 0 $\checkmark$ 0 $\checkmark$ 0 $\checkmark$ 0 $\checkmark$
AI 0 0 0 0 0 0 0 0 0
BDI $\checkmark$ 0 0 0 $\checkmark$ 0 $\checkmark$ $\checkmark$ $\checkmark$
D $\checkmark$ 0 0 0 0 0 0 $\checkmark$ $\checkmark$
DIII $\checkmark$ 0 $\checkmark 0 0$ 0 $\checkmark$ 0 $\checkmark$
AII 0 0 0 0 0 0 0 0 0
CII $\checkmark$ 0 $\checkmark$ $\checkmark$ $\checkmark$ 0 0 0 $\checkmark$
C 0 0 0 $\checkmark$ $\checkmark$ 0 0 0 0
CI 0 0 $\checkmark$ 0 $\checkmark$ 0 $\checkmark 0 0$
For the Wigner-Dyson classes A, AI, AII, which do not have a “special” energy $E=0$, the energy eigenvalues of a TUI may be chosen such that all eigenstates have negative energy. The resulting conventional invariants are always trivial, as indicated in Tab. 2. The opposite is true for chiral classes AIII, BDI and CII, where all conventional topological phases can be localized. As a consequence, the topology of, e.g., chiral TIs in $d=3$ does not imply the existence of delocalized bulk states Altland et al. (2024). Whether the superconducting classes C, CI, D, and DIII are Wannier localizable or not depends on the dimension and the symmetry class, as summarized in Table 2 (see SM for details). It has been shown recently Altland et al. (2024), that Wannier localizability implies detachability of the boundary modes from the bulk spectrum. Comparing the present results with the recent study that classifies detachable boundary modes Nakamura et al. (2025), we establish the equivalence of Wannier localizability and detachability of the boundary modes in conventional tenfold-way topological phases.
While topological equivalence as TUIs implies topological equivalence as conventional TI if $d>0$, the opposite is not true. In this sense, the TUI classification is more “refined” than the TI classification. A nontrivial example is the $\mathbb{Z}\times\mathbb{Z}$ classification of chiral TUIs with broken time-reversal symmetry for $d=3$ (see Table 1), which has to be contrasted with the $\mathbb{Z}$ classification of TIs: For each (integer-valued) conventional topological invariant, an additional integer-valued topological invariant is required to characterize the TUI. This additional invariant counts the number of finite-energy critical states appearing on the boundary SM ; Altland et al. (2024). Table 4 of SM gives an overview of this “beyond-tenfold-way” aspect of the TUI classification for all symmetry classes and all dimensions.
@@ D $\checkmark$ 0 0 0 0 0 0 $\checkmark$ $\checkmark$
Conclusions — Ultra-localized insulators, insulators that have exponentially localized states at all energies, can nevertheless be topological. The resulting topology is a global feature of the fully localized spectrum, and only transitions between topologically inequivalent ultra-insulators lead to metallic states somewhere in the spectrum. While these systems coincide with conventional topological matter in one dimension, and cannot exist in two dimensions, we uncovered many new instances of TUIs in three dimensions in different symmetry classes. As a hallmark their topology, TUIs host anomalous delocalized surface states that resist arbitrary surface disorder. Furthermore, TUIs provide new insights into TIs: by finding the ‘intersection’ of the TUI and TI classifications, we uncover many Wannierizable TIs.
@@
The fact that TUI and TI classifications differ is a consequence of the different objectives – studying equivalence classes of fully localized Hamiltonians versus those of ground states of gapped Hamiltonians. This change of viewpoint may be fruitful when applied to time-dependent quantum systems, such as Floquet and random unitary circuits, where local unitary operators (instead of Hamiltonians) play a central role Titum et al. (2016); Roy and Harper (2017); Long et al. (2021); Liu et al. (2023b); Fidkowski and Hastings (2023). It would then be desirable to extend our classification scheme of local unitaries Zhang and Levin (2021) beyond the single particle description.
@@
The construction of TUIs presented in this work corresponds to an infinite-disorder limit of a particular correlated disorder, instead of commonly considered independently distributed disorder. It is an open question whether some TUIs can be reached by adding generic (uncorrelated) disorder in a clean topological insulator. Nevertheless, the toy models presented in this work are potentially realizable in artificial platforms, such as electrical circuits Lee et al. (2018). Since the limit of infinite uncorrelated disorder gives rise to a trivial ultra-insulator, a competition between different disorders should result in a transition between TUIs, as sketched in Fig. 1(c). More extreme types of structural disorders, such as amorphous crystals, have recently been shown to realize conventional topological insulators Agarwala and Shenoy (2017); Mitchell et al. (2018); Marsal et al. (2020), but the full topological classification of disordered electronic matter is unknown, in particular extensions to disordered hyperuniform systems Torquato and Stillinger (2003); Zheng et al. (2020). Our work thus opens a pathway to generalize notions of topology to wider range of systems.
@@ Acknowledgements.
Acknowledgments — B.L. acknowledges financial support from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 214461). L.T. acknowledges support by the FNS/SNF Ambizione Grant No. PZ00P2_179962. T.N. acknowledges support from the Swiss National Science Foundation through a Consolidator Grant (iTQC, TMCG-2_2138052\_213805) P.B. is supported by the Deutsche Forschungsgemeinschaft (DFG) project Grant No. 277101999 within the CRC network TR 183 (subproject No. A03).
@@
[…]
@@
Correspondingly, the matrix $U$ is an orthogonal matrix and the auxiliary Hamiltonian $H_{\mathrm{aux}}$ is in tenfold-way class BDI, so that $K_{\rm aux}(d)={\cal K}_{\rm BDI}(d)$ for class AI. In this case, the “phase matrix” $D$ consists of signs $\pm 1$. Since relabeling matrices $P$ may have arbitrary integer winding number in $d=1$, but possess no nontrivial topology if $d>1$, it also follows that $K_{\rm aux}^{\prime}(d)={\cal K}_{\rm BDI}(1)=\mathbb{Z}$ if $d=1$ and $K_{\rm aux}^{\prime}(d)=0$ otherwise. We conclude that
@@
Class AII.— For class AII, $H$ must satisfy Eq. (4). Eigenvalues of $H$ come in twofold (Kramers) degenerate pairs. The same symmetry condition applies to the eigenvector matrix $U$, which makes it a symplectic matrix, and to the auxiliary Hamiltonian $H_{\mathrm{aux}}$, which places it in tenfold-way class CII. Since the relabeling matrix $P$ is proportional to the $2\times 2$ unit matrix $\sigma_{0}$, the corresponding winding numbers for $d=1$ are always even, so that $K_{\rm aux}^{\prime}(d)={\cal K}_{\rm CII}(1)=2\mathbb{Z}$ if $d=1$ and 0 otherwise. We then find
@@
where $\tau_{3}$ is the third Pauli matrix in a suitably chosen local basis. Further, in classes BDI and CII, $H$ is subject to an additional symmetry condition of the form $H=H^{*}$ or $H=\sigma_{2}H^{*}\sigma_{2}$. This additional symmetry condition commutes with $\tau_{3}$ and imposes that elements of $H$ are real or quaternion numbers. The antisymmetry condition (S8) also imposes that the eigenvalues come in pairs $\pm E_{\mathbf{n}}$ and that the corresponding eigenvectors are related to each other by multiplication by $\tau_{3}$. In class CII, all eigenvalues are twofold (Kramers) degenerate.
@@
where $U_{+}$ and $U_{-}$ are unitary, orthogonal, or symplectic matrices of equal size. The Hamiltonian $H$ then has the diagonal form (2) where $\Lambda$ is a block-diagonal matrix containing blocks $E_{\mathbf{n}}\tau_{1}$ for classes AIII and BDI and blocks $E_{\mathbf{n}}\tau_{1}\sigma_{0}$ for class CII, where each pair $\pm E_{\mathbf{n}}$ of eigenvalues of $H$ appears once. (The eigenvalues $E_{\mathbf{n}}$ appearing in $\Lambda$ may be positive or negative. Note that, strictly speaking, the matrix $U$ in this case cannot be called “eigenvector matrix”, since $\Lambda$ is not strictly diagonal.) Right-multiplication of $U=\mbox{diag}\,(U_{+},U_{-})$ with a relabeling matrix $P$ proportional to $\tau_{0}$ for classes AIII and BDI and to $\tau_{0}\sigma_{0}$ for class CII does not change $H$. Further, one may also right-multiply $U$ with a block-diagional “phase matrix” $D$. Its diagonal blocks $D_{\mathbf{n}}$ consist of $\tau_{0}$ or $\tau_{3}$ times a phase factor, sign $\pm 1$, or a $2\times 2$ matrix $\sigma\in\mbox{SU(2)}$ for classes AIII, BDI, and CII, respectively, while simultaneously changing the sign of the corresponding eigenvalue $E_{\mathbf{n}}$ in the eigenvalue matrix $\Lambda$ if $D_{\mathbf{n}}$ contains $\tau_{3}$.
@@
Its eigenvalues come in pairs $\pm E_{\mathbf{n}}$ and the eigenvectors are related to each other via particle-hole conjugation (i.e., multiplication with $\tau_{1}$ and complex conjugation). For an ultra-localized Hamiltonian we impose the additional condition that this also applies to the eigenvalue $E=0$, i.e., we rule out the possibility of a “topological zero-eigenvalue” for which the eigenstates are localized, but not related by particle-hole conjugation.
@@
where $\Lambda$ is a diagonal matrix containing blocks $E_{\mathbf{n}}\tau_{3}$, with each pair $\pm E_{\mathbf{n}}$ of eigenvalues of $H$ appearing once. The eigenvector matrix $U$ satisfies the antiunitary symmetry